A hard disk has average seek time 9ms, 7200 revolutions per minute, disk transfer rate
of 30 MBytes/sec and controller transfer rate of 500 MBits/sec. What is the average
time to read or write a 1KB sector to this hard disk?
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

Disk latency = Seek Time + Rotation Time + Transfer Time + Controller Overhead
Seek Time? Depends no. tracks the arm moves and seek speed of disk
Rotation Time? depends on rotational speed and how far the sector is from the head 
Transfer Time? depends on data rate (bandwidth) of disk (bit density) and the size of request

Disk latency = Seek Time + Rotation Time + 
                        Transfer Time + Controller Overhead

Average Rotational Time = (0.5)/(7200/60) = 4.166 miliseconds
[On average half rotation is made]

So Avg. Seek Time =  9 miliseconds.

Transfer Time = 1000/ (30 × 106 bytes/sec)
              = 33.33 microseconds

Controller Overhead = 500*125000/(30 × 106 bytes/sec)
                    = 0.208 miliseconds

Disk latency = Seek Time + Rotation Time + 
                           Transfer Time + Controller Overhead
             = 9 + 4.166 + 30.33 * 10-3 + 0.2 miliseconds
             = 13.4 miliseconds

Kindly revert for any queries

Thanks.